# READS-V: Real-time Automated Detection of Epileptic Seizures from Surveillance Videos via Skeleton-based Spatiotemporal ViG

### v1

@@ Abstract
An accurate and efficient epileptic seizure onset detection system can significantly benefit patients. Traditional diagnostic methods, primarily relying on electroencephalograms (EEGs), often result in cumbersome and non-portable solutions, making continuous patient monitoring challenging. The video-based seizure detection system is expected to free patients from the constraints of scalp or implanted EEG devices and enable remote monitoring in residential settings. Previous video-based methods neither enable all-day monitoring nor provide short detection latency due to insufficient resources and ineffective patient action recognition techniques. Additionally, skeleton-based action recognition approaches remain limitations in identifying subtle seizure-related actions. To address these challenges, we propose a novel skeleton-based spatiotemporal vision graph neural network (STViG) for efficient, accurate, and timely REal-time Automated Detection of epileptic Seizures from surveillance Videos (READS-V). Our experimental results indicate STViG outperforms previous state-of-the-art action recognition models on our collected patients’ video data with higher accuracy (5.9% error) and lower FLOPs (0.4G). Furthermore, by integrating a decision-making rule that combines output probabilities and an accumulative function, our READS-V system achieves a 5.1 s EEG onset detection latency, a 13.1 s advance in clinical onset detection, and zero false detection rate. The code is available at: https://github.com/xuyankun/STViG-for-READS-V
@@ 1 Introduction
As a common neurodegenerative disorder, epilepsy affects approximately $1\%$ population worldwide [27, 35]. A reliable epileptic seizure onset detection system can benefit patients significantly, because such a system equipped with accurate algorithms can promptly alert a seizure onset. According to previous studies [34, 32, 17], they predominantly focused on designing seizure detection algorithms based on electroencephalograms (EEGs). Although EEG can sense subtle brain changes just after a seizure begins, the use of scalp or implantable EEG devices often causes discomfort to patients and restricts them to hospital epilepsy monitoring units (EMUs). Consequently, there is a growing interest in developing an accurate video-based seizure detection system that could alleviate the discomfort associated with EEG helmets and facilitate remote monitoring of epileptic patients in residential settings [41]. However, the development of video-based seizure detection is hindered by several challenges as follows:
Figure 1: Motivation of skeleton-based patch embedding for seizure-related action recognition. The left shows real seizure-related actions; The middle shows challenges of traditional skeleton-based approaches; The right shows our strategy to address challenges.
@@
(1) Lack of datasets. Seizure-related video data collection is substantially time-consuming and requires doctoral expertise to annotate the different seizure-related periods. And, surveillance video recordings from the hospital EMUs contain highly sensitive information, they often capture the faces of patients, their families, and healthcare staff. This sensitivity raises privacy concerns and complicates the data collection process. Based on these reasons, there is no public video data intended for seizure study yet, which hinders the development of video-based seizure detection research.
@@
(2) Lack of effective analytic tools. Fundamentally, video-based seizure detection is an action recognition task, requiring the analysis of complex seizure-related actions in epileptic patients. Prior research indicates that existing methods struggle with achieving short detection latency and are often limited to nocturnal seizures when using raw RGB frames [46, 42]. While skeleton-based approaches powered by graph convolutional networks (GCNs) offer several advantages over RGB-based methods [9, 47, 21], there also encounter significant limitations in the context of seizure detection. Firstly, pre-trained pose estimation models [4, 10, 2, 26, 37] cannot be applied to epileptic patients’ action recognition directly due to the complexity of epileptic patient scenarios. These models suffer from unusual clothing, varied camera angles, and intricate behaviors of patients, leading to difficulties in accurately tracking patient skeletons. Secondly, as shown in Fig. 1, there are inherent difficulties in recognizing certain seizure-related actions using traditional skeleton-based approaches. For example, When a patient turns head, the tracking of the nose joint across frames can create ambiguity, making it unclear whether the patient is turning head or simply shifting it. Similarly, eye movements or wrist/ankle twists can only be tracked as static points instead of any changes. Furthermore, finger/toe movement also cannot be tracked by the model with the mainstream 17-keypoint template [24, 1].
@@
(3) Detection latency and false detection rate. Detection latency is a critical metric in a seizure detection system, representing the time gap between the real onset of a seizure and the detected onset. A shorter detection latency is highly desirable since it enables the system to alert caregivers before the onset of severe tonic-clonic symptoms, allowing timely interventions to prevent secondary injuries from a seizure attack. Also, false detection rate (FDR) is important in determining the quality of a seizure detection system [40, 36]. In terms of EEG time-series, the characteristics of EEG signals during the ictal period significantly differ from those in a healthy status, particularly in the frequency characteristics. In video-based detection, however, many normal patient behaviors closely resemble seizure-related actions. This similarity presents a challenge in accurately distinguishing between them, especially when decisions are based on a single video clip.
@@
In this study, we propose a novel skeleton-based spatiotemporal vision graph neural network (STViG), designed for efficient, accurate, and timely REal-time Automated Detection system of epileptic Seizures from surveillance Videos (READS-V). We address the aforementioned limitations by offering the following contributions: (1) We acquire a dataset of epileptic patient data from hospital EMUs and train a custom patient pose estimation model by manual pose annotations; (2) As expressed in the Fig. 1, instead of using skeleton-based coordinates, our STViG model utilizes skeleton-based patch embeddings as inputs to address remaining challenges. We also introduce a partitioning strategy in STViG to learn spatiotemporal inter-partition and intra-partition representations; (3) We generate the probabilities instead of binary classifications as outputs in this application, and make use of accumulative probability strategy to obtain shorter detection latency and lower FDR in real-time scenarios.
@@ 2 Related Work
Video-based seizure detection. There have been several efforts paid to video-based seizure detection or jerk detection since 2012 [6, 42, 19, 13], but they neither utilized DL approaches to achieve accurate performance nor supervised patients 24/7 by video monitoring. Since 2021, two recent works [46, 20] made use of DL models to improve the video-based seizure detection performance, however, one cannot provide all-day video monitoring with short latency and another one did not utilize a common RGB camera for video monitoring. Our proposed STViG framework can provide accurate and timely 24/7 video monitoring.
@@
Action recognition. There are two mainstream strategies for action recognition tasks, one is 3D-CNN for RGB-based action recognition, and the other one is skeleton-based action recognition. Many 3D-CNN architectures [12, 11, 18, 38, 39] have been proven to be effective tools for learning spatiotemporal representations of RGB-based video streams, they are widely applied to video understanding and action recognition tasks. Compared to RGB-based approaches with a large number of trainable parameters, skeleton-based approaches perform better in both accuracy and efficiency because they only focus on the skeleton information which is more relevant to the actions, and can alleviate contextual nuisances from raw RGB frames. Several GCN-based approaches [30, 23, 33, 5, 45] were proposed to achieve great performance in skeleton-based action recognition. [9] proposed PoseConv3D to combine 3D-CNN architecture and skeleton-based heatmaps, and proposed its variant RGBPoseConv3D to fuse RGB frames to obtain better performance. However, all aforementioned action recognition models show weaknesses in analyzing seizure-related actions, which include subtle behavioral changes when seizures begin. In this work, we propose STViG model to address this challenge.
Figure 2: Proposed skeleton-based STViG framework for READS-V system. Starting from raw RGB frames, we extract skeleton-based patches around each joint by fusing RGB frame and pose heatmap. These patches are then transformed into feature vectors through a patch embedding layer. We utilize a partitioning strategy to divide skeleton-based embeddings into several partitions. The STViG model, comprising both spatial and temporal modeling modules, is designed to learn inter- and intra-partition representations, thereby facilitating a better understanding of seizure-related actions. The bottom three figures express how the partitioning strategy works. After generating probability outputs, an accumulative probability strategy is used to enhance the performance of seizure onset decisions.
@@
Vision graph neural network. [14] first proposed vision graph neural networks (ViG) as an efficient and effective alternative backbone for image recognition. Since then many ViG variants [15, 28, 49, 43] have been proposed to handle various applications. Inspired by ViG, we first propose STViG to extend ViG to accomplish a skeleton-based action recognition task from videos.
3 READS-V Framework
@@
Fig. 2 presents an overview of skeleton-based STViG framework designed for READS-V. The process Starting begins with a video clip composed of consecutive RGB frames, we train the custom pose estimation model to generate joint heatmaps, then fuse these heatmaps and raw frames to construct skeleton-based patches. Prior to being processed by the STViG model, these patches are transformed into feature vectors through a trainable patch embedding layer. Additionally, the joints are grouped into several partitions, enhancing our model’s ability to recognize seizure-related actions. During the STViG processing phase, we conduct spatial and temporal modeling respectively, and utilize partition structure to learn inter-partition and intra-partition representations. STViG model outputs probabilities instead of binary classes for video clips, then a decision-making rule integrating probabilities and accumulative function can achieve shorter detection latency and lower FDR. The subsequent subsections will provide detailed explanations of each step in this process.
3.1 Custom pose estimation
@@
The STViG framework relies on a pose estimation algorithm to extract joint heatmaps at first. However, when we tested several mainstream pre-trained pose estimation models on our patient video data, the results were unsatisfactory. These models failed to accurately track the locations of patients and their joints. Therefore, we have to train our custom pose estimation model.
@@
We manually annotated 580 frames containing various behaviors during the both ictal periods and healthy status across all patients, then utilized lightweight-openpose [29] model as basis for our training for its efficient deployment. Eventually, we achieved a custom pose estimation model for all patients.
@@ 3.2 Patch extraction and patch embedding
In custom pose estimation algorithm, each joint is associated with a heatmap generated by a 2D Gaussian map [3]. We make use of this heatmap as filters to fuse the raw RGB frames, then extract a small patch around each joint from the fused image. However, determining the optimal size of the Gaussian maps to capture the regions of interest is a challenge during the training phase of the pose estimation model. To address this, we manually extract the patch $\mathbf{p}_{it}$ of $i$-th ($i=1,2...,N$) joint at frame $t$ by generating a Gaussian map with adjustable $\sigma$ for each joint to filter the raw RGB frame $\textbf{{I}}_{t}$:
$\mathbf{p}_{it}=exp(-\frac{(m-x_{it})^{2}+(n-y_{it})^{2}}{2\sigma^{2}})\otimes\textbf{{I}}_{t}(m,n)$ (1)
@@ 3.3 Graph construction with partition strategy
In this study, we proposed a partitioning strategy to construct graph edges between nodes. According to Openpose joint template, which contains 18 joints, we focus on 15 joints by excluding three (l_ear, r_ear, neck) as redundant. These 15 joints are divided into 5 partitions, each comprising three joints (head: nose, left/right eye; right arm: right wrist/elbow/shoulder; right leg: right hip/knee/ankle; left arm: left wrist/elbow/shoulder; left leg: left hip/knee/ankle). In our STViG, we conduct both spatial and temporal modeling for the graph. Spatial modeling involves constructing two subsets of edges based on different partition strategies: inter-partition edge set, denoted as $\mathcal{E}_{inter}=\left\{v_{it}v_{jt}|i\in\mathbf{Part}_{p1},j\in\mathbf{Part}_{p2},p1\neq p2\right\}$ and intra-partition edge set, denoted as $\mathcal{E}_{intra}=\left\{v_{it}v_{jt}|(i,j)\in\mathbf{Part}_{p}\right\}$, where $p,p1,p2=1,...,5$. As shown in the bottom left two schematic figures in Fig. 2, each node is connected to the nodes from all other partitions in the inter-partition step, and each node is only connected to the other nodes within the same partition in the intra-partition step. The bottom-left two figures of Fig. 2 visualize the graph construction with partitioning strategy.
For temporal modeling, we consider neighbors of each node within a $K_{S}\times K_{T}$ in both temporal and spatial dimensions. The temporal edge set is denoted as $\mathcal{E}_{T}=\left\{v_{it}v_{j(t+1)}|d(v_{it},v_{jt})\leq K_{S}/2,d(v_{it},v_{i(t+1)})\leq K_{T}/2\right\}$, where $d$ is the distance between two nodes. To facilitate the partitioning strategy, we arrange the joints with their partitions in a 1D sequence. Thus, if $K_{S}=3$, as shown in the bottom-right figure of Fig. 2, the middle joint of a partition aggregates only from intra-partition nodes over $K_{T}$ frames, while the border joint of a partition aggregates from inter-partition nodes over $K_{T}$ frames.
@@ 3.4 Partitioning spatiotemporal graph modeling
Spatial Modeling. The proposed skeleton-based STViG aims to process the input features $\mathbf{X}\in\mathbb{R}^{N\times T\times C}$ into the probabilities of seizure onset. Starting from input feature vectors, we first conduct spatial modeling to learn the spatial representations between nodes at each frame by inter-partition and intra-partition graph convolution operation as $\mathcal{G}^{{}^{\prime}}=F_{intra}(F_{inter}(\mathcal{G},\mathcal{W}_{inter}),\mathcal{W}_{intra})$, where $F$ is graph convolution operation, specifically we adopt max-relative graph convolution [22, 14] to aggregate and update the nodes with fully-connected (FC) layer and activation function:
@@
where $\mathcal{N}^{t}(\mathbf{x}_{it})$ is neighbor nodes of $\mathbf{x}_{it}$ at frame $t$, $Linear$ is a FC layer with learnable weights and $\sigma$ is activation function, e.g. ReLU. The only difference between $F_{inter}$ and $F_{intra}$ depends on $\mathcal{N}^{t}(\mathbf{x}_{it})$ is from $\mathcal{E}_{inter}$ or $\mathcal{E}_{intra}$. Given input feature vectors at frame $t$, denoted as $\mathbf{X}_{t}\in\mathbb{R}^{N\times C}$, the graph convolution processing as Eq. 2 can be denoted as $\mathbf{X}^{{}^{\prime}}_{t}=MRGC(\mathbf{X}_{t})$, then the partitioning spatial graph processing module in STViG would be:
$\mathbf{X}_{t}^{{}^{\prime\prime}}$ $=\sigma(MRGC(\mathbf{X}_{t}W_{in\_inter})W_{out\_inter}+X_{t})$ (3)
$\mathbf{X}_{t}^{{}^{\prime}}$ $=\sigma(MRGC(\mathbf{X}_{t}^{{}^{\prime\prime}}W_{in\_intra})W_{out\_intra}+X_{t})$

### v2

# VSViG: Real-time Video-based Seizure Detection via Skeleton-based Spatiotemporal ViG

@@ Abstract
An accurate and efficient epileptic seizure onset detection can significantly benefit patients. Traditional diagnostic methods, primarily relying on electroencephalograms (EEGs), often result in cumbersome and non-portable solutions, making continuous patient monitoring challenging. The video-based seizure detection system is expected to free patients from the constraints of scalp or implanted EEG devices and enable remote monitoring in residential settings. Previous video-based methods neither enable all-day monitoring nor provide short detection latency due to insufficient resources and ineffective patient action recognition techniques. Additionally, skeleton-based action recognition approaches remain limitations in identifying subtle seizure-related actions. To address these challenges, we propose a novel Video-based Seizure detection model via a skeleton-based spatiotemporal Vision Graph neural network (VSViG) for its efficient, accurate and timely purpose in real-time scenarios. Our experimental results indicate VSViG outperforms previous state-of-the-art action recognition models on our collected patients’ video data with higher accuracy (5.9% error), lower FLOPs (0.4G), and smaller model size (1.4M). Furthermore, by integrating a decision-making rule that combines output probabilities and an accumulative function, we achieve a 5.1 s detection latency after EEG onset, a 13.1 s detection advance before clinical onset, and a zero false detection rate. The project homepage is available at: https://github.com/xuyankun/VSViG/
Keywords:
Video-based seizure detection Action recognition Spatiotemporal vision GNN
@@ 1 Introduction
As a common neurodegenerative disorder, epilepsy affects approximately $1\%$ population worldwide [31, 40]. A reliable epileptic seizure onset detection method can benefit patients significantly, because such a system equipped with accurate algorithms can promptly alert a seizure onset. According to previous studies [39, 37, 19], they predominantly focused on designing seizure detection algorithms based on electroencephalograms (EEGs). Although EEG can sense subtle brain changes just after a seizure begins, the use of scalp or implantable EEG devices often causes discomfort to patients and restricts them to hospital epilepsy monitoring units (EMUs). Consequently, there is a growing interest in developing an accurate video-based seizure detection system that could alleviate the discomfort associated with EEG helmets and facilitate remote monitoring of epileptic patients in residential settings [25]. However, the development of video-based seizure detection is hindered by several challenges as follows:
@@
(1) Lack of datasets. Seizure-related video data collection is substantially time-consuming and requires doctoral expertise to annotate the different seizure-related periods. And, surveillance video recordings from the hospital EMUs contain highly sensitive information. Based on these reasons, there is no public video data intended for seizure study yet. Although [34] released a related dataset, they only implement action anticipation on ictal videos which cannot provide any false detection rate (FDR) information.
Figure 1: Motivation of proposed skeleton-based patch embedding. The left shows real seizure-related actions; The middle shows challenges of traditional skeleton-based approaches; The right shows our strategy to address challenges.
Figure 2: Model comparison of achieved errors and the number of parameters.
@@
(2) Lack of effective analytic tools. Fundamentally, video-based seizure detection is an action recognition task, requiring the analysis of complex seizure-related actions in epileptic patients. Prior research indicates that existing methods struggle with achieving short detection latency and are often limited to nocturnal seizures when using raw RGB frames [51, 46]. While skeleton-based approaches powered by graph convolutional networks (GCNs) offer several advantages over RGB-based methods [11, 52, 23], there also encounter significant limitations in the context of seizure detection. Firstly, pre-trained pose estimation models [4, 12, 2, 28, 42] cannot be applied to epileptic patients’ action recognition directly due to the complexity of epileptic patient scenarios, as shown in Fig. 6. Secondly, as shown in Fig. 1, there are inherent difficulties in recognizing certain seizure-related actions naively using coordinates as traditional skeleton-based approaches did.
@@
(3) Detection latency and false detection rate. Detection latency is a critical metric in a seizure detection system, representing the time gap between the real onset of a seizure and the detected onset. A shorter detection latency, known as early seizure detection is highly desirable since it enables timely interventions prior to a serious seizure attack. Also, FDR is important [45, 41], however, in video-based detection many normal patient behaviors closely resemble seizure-related actions. Previous studies are hard to distinguish them or even directly overlook this metric.
@@
In this study, we propose a novel Video-based Seizure detection model via skeleton-based spatiotemporal Vision Graph neural network (VSViG) intended for efficient, accurate, and timely real-time scenarios. We address the aforementioned limitations by offering the following contributions: (1) We acquire a dataset of epileptic patient data from hospital EMUs and fine-tune the pose estimation model for epileptic patients; (2) As expressed in the Fig. 1, instead of using skeleton-based coordinates, our VSViG model utilizes skeleton-based patch embeddings as inputs to address mentioned challenges. Our proposed model including base and light versions outperforms the state-of-the-art models from the field of action recognition and video-based seizure detection in both accuracy and model size, as shown in Fig. 2; (3) We define the application as a regression task to generate the probability/likelihood outputs instead of naive binary classification on interictal (healthy) and ictal (unhealthy) status, thereby the model can accurately detect subtle or trending seizure-related actions in low likelihood for early detection propose.
@@ 2 Related Work
Video-based seizure detection. There have been several efforts paid to video-based seizure detection or jerk detection over the past decade [8, 46, 21, 15], but they neither achieved accurate performance nor supervised patients 24/7 by video monitoring. Recently, several works [51, 22, 30, 34] made use of advanced deep learning (DL) models to improve the video-based seizure detection performance, however, there are many limitations among these works: (1) [51, 22] cannot achieve short detection latency or require more video modality data than RGB; (2) [30, 34] defined the early seizure detection task as a video-based action anticipation task, but they cannot achieve higher accuracy until utilizing at least 1/2 clip of ictal videos which cannot bring short latency, further action anticipation task cannot provide any FDR information.
@@
Action recognition. There are two mainstream strategies for action recognition tasks, one is 3D-CNN for RGB-based action recognition, and the other one is skeleton-based action recognition. Many 3D-CNN or ViT architectures [14, 13, 20, 43, 44, 6, 48, 1] have been proven to be effective tools for learning spatiotemporal representations of RGB-based video streams, they are widely applied to video understanding and action recognition tasks. Compared to RGB-based approaches with a large number of trainable parameters, skeleton-based approaches perform better in both accuracy and efficiency because they only focus on the skeleton information which is more relevant to the actions, and can alleviate contextual nuisances from raw RGB frames. Several GCN-based approaches [35, 27, 38, 7, 50] were proposed to achieve great performance in skeleton-based action recognition. [11] proposed PoseConv3D to combine 3D-CNN architecture and skeleton-based heatmaps, and proposed its variant RGBPoseConv3D to fuse RGB frames to obtain better performance. However, all aforementioned action recognition models show weaknesses in analyzing seizure-related actions, which include subtle behavioral changes when seizures begin. In this work, we propose VSViG model to address this challenge.
@@
Vision graph neural network. [16] first proposed vision graph neural networks (ViG) as an efficient and effective alternative backbone for image recognition. Since then many ViG variants [17, 32, 54, 47] have been proposed to handle various applications. Inspired by ViG, we first propose VSViG to extend ViG to accomplish a skeleton-based action recognition task from seizure-related videos.
Figure 3: Proposed skeleton-based VSViG framework. Starting from raw RGB frames, we extract skeleton-based patches around each joint by fusing RGB frames and pose heatmaps. Then features from patches are generated by a patch embedding. In spatiotemporal ViG modeling, a partition strategy with proposed inter-, intra-, and dynamic partition operations is used, as shown in the bottom three subfigures.
3 VSViG Framework
@@
Fig. 3 presents an overview of skeleton-based VSViG framework. The process begins with a video clip composed of consecutive RGB frames, we fine-tune the pose estimation model to generate joint heatmaps, then fuse these heatmaps and raw frames to construct skeleton-based patches. Before being processed by the VSViG model, these patches are transformed into feature vectors through a trainable patch embedding layer. During the spatiotemporal ViG processing phase, we conduct spatial and temporal modeling respectively, and utilize partition structure to learn inter-partition and intra-partition representations. VSViG model outputs probabilities instead of binary classes for video clips, then a decision-making rule integrating probabilities and accumulative function can achieve shorter detection latency and lower FDR. The subsequent subsections will provide detailed explanations of each step in this process.
3.1 Pose estimation model fine-tuning
@@
The VSViG framework relies on a pose estimation algorithm to extract joint heatmaps at first. However, when we tested several mainstream pre-trained pose estimation models on our patient video data, the results were unsatisfactory, which are shown as Fig. 6. These models failed to accurately track the locations of patients and their joints. Therefore, we have to fine-tune the pose estimation model for epileptic patients.
@@
We manually annotated 580 frames containing various behaviors during both ictal periods and healthy status across all patients, then utilized lightweight-openpose [33] as a base model to fine-tune due to its efficiency. Eventually, we achieved a fine-tuned pose estimation model can accurately track patients’ joints across all patients, as shown in Fig. 6.
@@ 3.2 Patch extraction and patch embedding
In custom pose estimation algorithm, each joint is associated with a heatmap generated by a 2D Gaussian map [5]. We make use of this heatmap as a filter to fuse the raw RGB frames, then extract a small patch around each joint from the fused image. However, determining the optimal size of the Gaussian maps to capture the regions of interest is a challenge during the training phase of the pose estimation model. To address this, we manually extract the patch $\mathbf{p}_{it}$ of $i$-th ($i=1,2...,N$) joint at frame $t$ by generating a Gaussian map with adjustable $\sigma$ for each joint to filter the raw RGB frame $\textbf{{I}}_{t}$:
$\mathbf{p}_{it}=\text{exp}(-\frac{(m-x_{it})^{2}+(n-y_{it})^{2}}{2\sigma^{2}})\otimes\textbf{{I}}_{t}(m,n)$ (1)
@@ 3.3 Graph construction with partition strategy
In this study, we proposed a partitioning strategy to construct graph edges between nodes. According to Openpose joint template, which contains 18 joints, we focus on 15 joints by excluding three (l_ear, r_ear, neck) as redundant. These 15 joints are divided into 5 partitions, each comprising three joints (head: nose, left/right eye; right arm: right wrist/elbow/shoulder; right leg: right hip/knee/ankle; left arm: left wrist/elbow/shoulder; left leg: left hip/knee/ankle). In our VSViG, we conduct both spatial and temporal modeling for the graph. Spatial modeling involves constructing two subsets of edges based on different partition strategies: inter-partition edge set and intra-partition edge set, which are denoted as:
$\mathcal{E}_{inter}$ =$=\left\{v_{it}v_{jt}|i\in\mathbf{Part}_{p_{1}},j\in\mathbf{Part}_{p_{2}},p_{1}\neq p_{2}\right\}$ (2)
$\mathcal{E}_{intra}$ =$=\left\{v_{it}v_{jt}|(i,j)\in\mathbf{Part}_{p}\right\}$
where $p/p_{1}/p_{2}=1,...,5$. As shown in the bottom left two schematic figures in Fig. 3, each node is connected to the nodes from all other partitions in the inter-partition step, and each node is only connected to the other nodes within the same partition in the intra-partition step. The bottom-left two figures of Fig. 3 visualize the graph construction with partitioning strategy.
For temporal modeling, we consider neighbors of each node within a $K_{S}\times K_{T}$ in both temporal and spatial dimensions. The temporal edge set is denoted as $\mathcal{E}_{T}=\left\{v_{it}v_{j(t+1)}|d(v_{it},v_{jt})\leq K_{S}/2,d(v_{it},v_{i(t+1)})\leq K_{T}/2\right\}$, where $d$ is the distance between two nodes. To facilitate the partitioning strategy, we arrange the joints with their partitions in a 1D sequence. Thus, if $K_{S}=3$, as shown in the bottom-right of Fig. 3, the middle joint of a partition aggregates only from intra-partition nodes over $K_{T}$ frames, while the border joint of a partition aggregates from inter-partition nodes over $K_{T}$ frames.
@@ 3.4 Partitioning spatiotemporal graph modeling
Spatial Modeling. The proposed skeleton-based VSViG aims to process the input features $\mathbf{X}\in\mathbb{R}^{N\times T\times C}$ into the probabilities of seizure onset. Starting from input feature vectors, we first conduct spatial modeling to learn the spatial representations between nodes at each frame by inter-partition and intra-partition graph convolution operation as $\mathcal{G}^{{}^{\prime}}=F_{intra}(F_{inter}(\mathcal{G},\mathcal{W}_{inter}),\mathcal{W}_{intra})$, where $F$ is graph convolution operation, specifically we adopt max-relative graph convolution [26, 16] to aggregate and update the nodes with fully-connected (FC) layer and activation function:
@@
where $\mathcal{N}^{t}(\mathbf{x}_{it})$ is neighbor nodes of $\mathbf{x}_{it}$ at frame $t$, $Linear$ is a FC layer with learnable weights and $\sigma$ is activation function, e.g. ReLU. The only difference between $F_{inter}$ and $F_{intra}$ depends on $\mathcal{N}^{t}(\mathbf{x}_{it})$ is from $\mathcal{E}_{inter}$ or $\mathcal{E}_{intra}$. Given input feature vectors at frame $t$, denoted as $\mathbf{X}_{t}\in\mathbb{R}^{N\times C}$, the graph convolution processing as Eq. 3 can be denoted as $\mathbf{X}^{{}^{\prime}}_{t}=MRGC(\mathbf{X}_{t})$, then the partitioning spatial graph processing module in VSViG would be:
$\mathbf{X}_{t}^{{}^{\prime\prime}}$ $=\sigma(MRGC(\mathbf{X}_{t}W_{in\_inter})W_{out\_inter}+\mathbf{X})$ (4)
$\mathbf{X}_{t}^{{}^{\prime}}$ $=\sigma(MRGC(\mathbf{X}_{t}^{{}^{\prime\prime}}W_{in\_intra})W_{out\_intra}+\mathbf{X})$